# Hidden zero modes and topology of multiband non-Hermitian systems

### v2

@@ Abstract
In a finite one-dimensional non-Hermitian system, the number of zero modes does not necessarily reflect the topology of the system. This is known as the breakdown of the bulk-boundary correspondence and has lead to misconceptions about the topological protection of edge modes in such systems. Here we show why this breakdown does occur and that it typically results in hidden zero modes, extremely long-lived zero energy excitations, which are only revealed when considering the singular value instead of the eigenvalue spectrum. We point out, furthermore, that in a finite multiband non-Hermitian system with Hamiltonian $H$, one needs to consider also the reflected Hamiltonian $\tilde{H}$, which is in general distinct from the adjoint $H^{\dagger}$, to properly relate the number of protected zeroes to the winding number of $H$.
@@ I Introduction
In quantum physics, we usually consider observables which are Hermitian operators. These operators have a real eigenspectrum, guaranteeing that expectation values are real and time evolution unitary. However, it has been pointed out about 20 years ago that one can replace the condition of Hermiticity by the less stringent condition of space-time reflection (PT) symmetry and, if this symmetry is unbroken, still obtain a real spectrum [1]. Another motivation to study non-Hermitian Hamiltonians comes from Master equations for open quantum systems [2, 3]. If one ignores quantum jumps, which is a reasonable approximation in certain limits, one can rewrite such a Master equation as a non-Hermitian Hamiltonian [4, 5]. This approach has been used in recent times to analyze and interpret experimental results in optical and magnetic systems [6, 7, 8, 9].
Non-Hermitian systems do show a number of phenomena which are not present in the Hermitian case. Most intriguingly, their spectrum is extremely sensitive to small perturbations [10], see also App. A. The best known example is the non-Hermitian skin effect: changing the boundary conditions from periodic to open can lead to a localization of a macroscopic number of states at the boundaries [11]. Another well known phenomenon are exceptional points. At these points, two eigenvalues and also their corresponding eigenvectors coalesce [12, 13]. This is, of course, not possible in a Hermitian system where the eigenvectors always form an orthogonal system.
@@
From a theoretical perspective, an important step to understand these phenomena is to classify Gaussian non-Hermitian systems [14, 15]. Here, topology plays a crucial role because it is robust against small perturbations. In one-dimensional Hermitian systems, topological order is only possible if the system possesses additional symmetries. In the case of non-spatial symmetries—such as time reversal, particle-hole, and chiral symmetry—this leads to the tenfold classification of symmetry-protected topological (SPT) order [16]. In contrast, one-dimensional non-Hermitian systems can have a non-trivial topology even without additional symmetries because the eigenspectrum is complex and the determinant of the Bloch Hamiltonian $\det h(k):\,[0,2\pi)\to\mathbb{C}$ as function of momentum $k$ can have a non-zero winding number in the complex plane. It has been shown that if such a winding around a reference point $E$ in the complex plane exists for a single-band model then there will be a skin effect for open boundaries [11]. It has therefore been suggested that the standard bulk topological invariants are not useful to define a bulk-boundary correspondence for non-Hermitian systems and that they have to be replaced by other invariants [17, 18] or by an invariant based on a modified Hermitian Bloch Hamiltonian [15]. However, zero energy edge modes in finite non-Hermitian systems are, in general, fragile to perturbations [10, 19] (see also App. A), implying that they often lack topological protection thus making a bulk-boundary correspondence based on such modes questionable. As an alternative, the singular value spectrum has been put forward [20, 21, 22, 23]. The singular values $s_{i}$ of a system with Hamiltonian $H$ are the square roots of the eigenvalues of $H^{\dagger}H$. For a Hermitian system, we therefore have that $s_{i}=|\lambda_{i}|$ where $|\lambda_{i}|$ are the eigenvalues of $H$. In this case, the standard bulk-boundary correspondence also applies to the singular value spectrum. However, for a non-Hermitian $H$ the eigenvalue and the singular value spectrum are different. While the former is unstable to small perturbations, the latter is stable and important properties can be directly inferred from the topological winding number.
@@
In this article we will show why this is the case. We show, in particular, that non-Hermitian systems quite generally have hidden zero modes if they have a non-trivial topology. These are topologically protected edge modes with exact zero eigenvalues for semi-infinite boundaries which are, however, not present and do not converge to zero with system size for a finite system. They do, however, get mapped exponentially close to zero by a finite-system Hamiltonian and thus represent extremely long-lived states which are physically highly relevant and might be experimentally indistinguishable from true eigenstates. We will show, furthermore, that in the multiband case one has to consider not only the Hamiltonian $H$ but also the reflected Hamiltonian $\tilde{H}$. We will put the relation between topology, zero modes for semi-infinite boundaries, and protected singular values for finite systems which converge to zero with increasing system size, on a firm footing by using theorems known from the study of Toeplitz operators. Using several examples, we will highlight that the eigenspectrum is, in general, insensitive to the topology of the system while the singular value spectrum can be directly related to the winding number. In App. A we will show, furthermore, that the bulk-boundary correspondence put forward here fully explains the topological protection of stable edge modes in models considered previously in the literature. On the other hand, the approaches considered in Refs. [17, 18, 15] do predict zero modes in open systems, however, these modes generically are, as we will show, unstable to small perturbations and thus not topologically protected.
@@ II Topology and winding numbers
In the second line, we have assumed that $h(k)$ is diagonalizable with eigenvalues $\lambda_{j}(k)$, $j=1,\cdots,N$. Note that $\lambda_{j}(0)=\lambda_{j}(2\pi)$ is periodic. This implies, in particular, that if all eigenvalues are real then $\mathcal{I}=0$. A generic one-dimensional Hermitian system, i.e. a system without additional symmetries, is thus always topologically trivial. The only way for a one-dimensional Hermitian system to have non-trivial topology is symmetry-protected topological (SPT) order. These symmetries lead to a block structure of $h(k)$ and topology can be defined based on the properties of these blocks. This leads to the tenfold classification scheme for non-spatial symmetries and to topological crystalline orders for spatial symmetries [24, 25, 26, 27]. In contrast, $\mathcal{I}\neq 0$ is a generic property of non-Hermitian systems which does not require the presence of additional symmetries. The bulk-boundary correspondence for SPT phases of Hermitian systems connects the bulk topological invariant with the number of gapless protected edge modes in a system with boundaries. In a non-Hermitian system, on the other hand, which has a non-zero winding number around a reference energy $E$, we have the skin effect [11]. I.e., a macroscopic number of modes become localized in a system with boundaries but they are fragile with respect to small perturbations.
@@ III Topology in one-dimensional semi-infinite systems
Thus, the winding number $\mathcal{I}$ tells us directly about the difference in the number of right and left zero eigenvalues of $H$. It is also very important to distinguish between the case where $h_{j}$ are $N\times N$ blocks with $N>1$ and the scalar case, $N=1$, where $h_{j}$ are just complex numbers. In the latter case, Coburn’s lemma states that either $D(\text{ker}(H))=0$ or $D(\text{ker}(H^{\dagger}))=0$. I.e., in this case the sign of a non-zero winding number $\mathcal{I}$ tells us immediately whether $H$ has right or left zeroes and $|\mathcal{I}|$ gives the number of such zeroes.
III.1 Example 1
To illustrate these points, we consider an example where $h(k)=1+2\text{e}^{-\text{i}k}$ which means that $\mathcal{I}=-\text{ind}(H)=-1$ and the Fourier coefficients are $h_{0}=1$ and $h_{1}=2$ with all the other ones being equal to zero. Coburn’s lemma then implies that $D(\text{ker}(H))=1$ and $D(\text{ker}(H^{\dagger}))=0$, i.e., $H$ has exactly one right zero mode. In this simple case, we can explicitly construct the right zero mode by considering $Hv=0$ with $v=\begin{pmatrix}v_{1}&v_{2}&\dots\end{pmatrix}^{T}$ and $v_{j}\in\mathbb{C}$. This leads to the recurrence relation for the vector coefficients $v_{j}+2v_{j+1}=0$. A normalized solution is given by $v_{j}=(-1)^{j-1}\sqrt{3}/2^{j}$ which means that the zero mode is exponentially localized.
@@ IV Finite systems and hidden zero modes
In a finite Hermitian system with non-trivial SPT order and system size $L$, zero modes can typically be easily identified as eigenvalues which scale as $|\lambda|\sim\exp(-L)$. This, however, is in general not the case in a non-Hermitian system. Here, zero modes can be hidden. This can be understood by considering the example discussed above. If we consider $H_{L}v=0$ for a finite matrix $H_{L}$ and vector $v=\begin{pmatrix}v_{1}&v_{2}&\dots&v_{L}\end{pmatrix}^{T}$ then the recurrence relation above is supplemented by the boundary condition $v_{L}=0$. In this case, the only solution is $v\equiv 0$ which is not a proper solution. The protected zero mode, which does exist in the thermodynamic limit, cannot be found by considering the eigenspectrum of a finite system. The only eigenstate for a finite system is $v=\begin{pmatrix}1&0&\dots 0\end{pmatrix}^{T}$ with eigenvalue $E=1$.
@@
In the example, the vector $v=\begin{pmatrix}v_{1}&v_{2}&\dots&v_{L}\end{pmatrix}^{T}$ with $v_{j}=(-1)^{j-1}\sqrt{3}/2^{j}$ for $j<L$ and $v_{L}=0$ is a hidden zero mode with $\lim_{L\to\infty}||H_{L}v||=\lim_{L\to\infty}|v_{L}|=0$.
@@
IV.1 Example 2
This is best illustrated by another example. Consider the Hamiltonian in $k$-space given by
@@ IV.1 Example 2
Figure 1: The Hamiltonian $H$ in Eq. (7), its reflection $\tilde{H}$, and the adjoint operators. The arrows indicate unidirectional hopping and the open circles the localized edge modes.
If we consider this Hamiltonian on a finite line then we find two exact zero modes which are localized at the right boundary. However, adding a small perturbation one zero mode can be removed and the other one becomes a hidden zero mode.
@@ V Singular values and hidden zero modes
The block case, $N>1$, is slightly more complicated. We know that $\tilde{\mathcal{I}}=-\mathcal{I}$ which implies that we can also write $K=D(\mbox{ker}(H^{\dagger}))+D(\mbox{ker}(\tilde{H}^{\dagger}))$ and therefore
@@
I.e., in the block case we only know that there are at least $|\mathcal{I}|\leq K$ singular values which will go to zero. Nevertheless, we obtain in both cases a clear bulk-boundary correspondence between the bulk winding number and the singular value spectrum of a finite system.
@@
where we have used that $U$ is unitary. We thus find one of the main results of this letter: Every singular value $s_{n}$ which vanishes in the thermodynamic limit is directly connected to a (hidden) zero mode $v_{n}$ of the Hamiltonian $H_{L}$. The $K$-splitting theorem then connects this to the winding number. For a system with winding number $\mathcal{I}$ there are at least $|\mathcal{I}|$ exact or hidden zero modes. In the scalar case $N=1$, there are exactly $|\mathcal{I}|$ many.
@@ VI Hidden zero modes in a sub-lattice symmetric multiband model
The case $\alpha=-1$ further exemplifies some of the findings in this work. Now $\mathcal{I}_{2}=-1$ if $x>0.9y$ and $\mathcal{I}_{2}=0$ otherwise meaning that the total winding $\mathcal{I}$ is either zero or $1$. Naively, we might expect that for $\mathcal{I}=0$ the topology is trivial and there is no skin effect. This, however, is incorrect. With sub-lattice symmetry being present, there are two topological invariants which can be either considered to be the windings of the blocks $\mathcal{I}_{1,2}$ or the sum and the difference of these two invariants [15, 31, 32]. The number of singular values $K$ is bounded from below by $K\geq|\mathcal{I}_{1}|+|\mathcal{I}_{2}|$. We thus expect to have either two or just one protected singular value consistent with the numerical results shown in Fig. 3. However, the single zero mode for the $\mathcal{I}_{1}=1$, $\mathcal{I}_{2}=0$ case is hidden while two eigenvalues which go to zero are present for the $\mathcal{I}_{1}=1$, $\mathcal{I}_{2}=-1$ case. The latter result can be explained by noting that the system for $\mathcal{I}_{1}=-\mathcal{I}_{2}$ can be adiabatically deformed to a Hermitian system. The two zero eigenmodes then follow from the standard bulk-boundary correspondence, see also App. A.

### v3

@@ Abstract
In a finite one-dimensional non-Hermitian system, the number of zero modes does not necessarily reflect the topology of the system. This is known as the breakdown of the bulk-boundary correspondence and has led to misconceptions about the topological protection of edge modes in such systems. Here we show why this breakdown does occur and that it typically results in hidden zero modes, extremely long-lived zero energy excitations, which are only revealed when considering the singular value instead of the eigenvalue spectrum. We point out, furthermore, that in a finite multiband non-Hermitian system with Hamiltonian $H$, one needs to consider also the reflected Hamiltonian $\tilde{H}$, which is in general distinct from the adjoint $H^{\dagger}$, to properly relate the number of protected zeroes to the winding number of $H$.
@@ I Introduction
In quantum physics, we usually consider observables which are Hermitian operators. These operators have a real eigenspectrum, guaranteeing that expectation values are real and time evolution unitary. However, it has been pointed out about 20 years ago that one can replace the condition of Hermiticity by the less stringent condition of space-time reflection (PT) symmetry and, if this symmetry is unbroken, still obtain a real spectrum [1]. Another motivation to study non-Hermitian Hamiltonians comes from Master equations for open quantum systems [2, 3]. If one ignores quantum jumps, which is a reasonable approximation in certain limits, one can rewrite such a Master equation as a non-Hermitian Hamiltonian [4, 5]. This approach has been used in recent times to analyze and interpret experimental results in optical and magnetic systems [6, 7, 8, 9]. Non-Hermitian systems do show a number of phenomena which are not present in the Hermitian case. Most intriguingly, their eigenspectrum is extremely sensitive to small perturbations [10]. The best known example is the non-Hermitian skin effect: changing the boundary conditions from periodic (PBC) to open (OBC) can lead to a localization of a macroscopic number of states at the boundaries [11]. Another well known phenomenon are exceptional points. At these points, two eigenvalues and also their corresponding eigenvectors coalesce [12, 13]. This is, of course, not possible in a Hermitian system where the eigenvectors always form an orthogonal system.
@@
From a theoretical perspective, an important step to understand these phenomena is to classify Gaussian non-Hermitian systems [14, 15]. Here, topology plays a crucial role because it is robust against small perturbations which break translational invariance. In one-dimensional Hermitian systems, topological order is only possible if the system possesses additional symmetries. In the case of non-spatial symmetries—such as time reversal, particle-hole, and chiral symmetry—this leads to the tenfold classification of symmetry-protected topological (SPT) order [16]. In contrast, one-dimensional non-Hermitian systems can have a non-trivial topology even without additional symmetries because the eigenspectrum is complex and the determinant of the Bloch Hamiltonian $\det h(k):\,[0,2\pi)\to\mathbb{C}$ as function of momentum $k$ can have a non-zero winding number in the complex plane. It has been shown that if such a winding around a reference point $E$ in the complex plane exists for a single-band model then there will be a skin effect for open boundaries [11]. It has therefore been suggested that the standard bulk topological invariants are not useful to define a bulk-boundary correspondence for non-Hermitian systems and that they have to be replaced by other invariants [17, 18] or by an invariant based on a modified Hermitian Bloch Hamiltonian [15]. However, zero energy edge modes in finite non-Hermitian systems are, in general, fragile to perturbations [10, 19], implying that they often lack topological protection thus making a bulk-boundary correspondence based on such modes questionable. It is crucial to note that "topological features must be properties of a translationally not invariant Hamiltonian" [20]. This is discussed further in App. A. As an alternative, the singular value spectrum has been put forward [21, 22, 23, 24]. The singular values $s_{i}$ of a system with Hamiltonian $H$ are the square roots of the eigenvalues of $H^{\dagger}H$. For a Hermitian system, we therefore have that $s_{i}=|\lambda_{i}|$ where $|\lambda_{i}|$ are the eigenvalues of $H$. In this case, the standard bulk-boundary correspondence also applies to the singular value spectrum. However, for a non-Hermitian $H$ the eigenvalue and the singular value spectrum are different. While the former is unstable to small perturbations, the latter is stable and important properties of the singular value spectrum can be directly inferred from the topological winding number.
@@
In this article we will show why this is the case. We show, in particular, that non-Hermitian systems quite generally have hidden zero modes if they have a non-trivial topology. These are topologically protected edge modes with exact zero eigenvalues for semi-infinite boundaries which are, however, not present in the eigenspectrum and do not converge to zero with system size for a finite system. They do, however, get mapped exponentially close to zero by a finite-system Hamiltonian and thus represent extremely long-lived states which are physically highly relevant and are only distinguishable experimentally from true eigenstates in the dynamics at a time scale which increases linearly with system size. We will show, furthermore, that in the multiband case one has to consider not only the Hamiltonian $H$ but also the reflected Hamiltonian $\tilde{H}$. We will put the relation between topology, zero modes for semi-infinite boundaries, and protected singular values for finite systems which converge to zero with increasing system size, on a firm footing by using theorems known from the study of Toeplitz operators. Using several examples, we will highlight that the eigenspectrum is, in general, insensitive to the topology of the system while the singular value spectrum can be directly related to the winding number. In App. A we will show, furthermore, that the bulk-boundary correspondence put forward here fully explains the topological protection of stable edge modes in models considered previously in the literature. On the other hand, the approaches considered in Refs. [17, 18, 15] do predict zero modes in open systems, however, these modes generically are, as we will show, related to translational invariance, unstable to small perturbations, and thus not topologically protected.
@@ II Topology and winding numbers
In the second line, we have assumed that $h(k)$ is diagonalizable with eigenvalues $\lambda_{j}(k)$, $j=1,\cdots,N$. Note that $\lambda_{j}(0)=\lambda_{j}(2\pi)$ is periodic. This implies, in particular, that if all eigenvalues are real then $\mathcal{I}=0$. A generic one-dimensional Hermitian system, i.e. a system without additional symmetries, is thus always topologically trivial. The only way for a one-dimensional Hermitian system to have non-trivial topology is symmetry-protected topological (SPT) order. These symmetries lead to a block structure of $h(k)$ and topology can be defined based on the properties of these blocks. This leads to the tenfold classification scheme for non-spatial symmetries and to topological crystalline orders for spatial symmetries [25, 26, 27, 28]. In contrast, $\mathcal{I}\neq 0$ is a generic property of non-Hermitian systems which does not require the presence of additional symmetries. The bulk-boundary correspondence for SPT phases of Hermitian systems uses the bulk topological invariant to provide a lower bound on the number of protected edge modes in a system with boundaries. In a non-Hermitian system, on the other hand, which has a non-zero winding number around a reference energy $E$, we have the skin effect [11]. I.e., a macroscopic number of modes become localized in a system with boundaries but they are fragile with respect to small perturbations.
@@ III Topology in one-dimensional semi-infinite systems
Thus, the winding number $\mathcal{I}$ tells us directly about the difference in the number of right and left zero eigenvalues of $H$ but not their total numbers. This is similar to the Chern number in the Hermitian case which describes the difference between left and right propagating edge modes but does not fix their total numbers.
It is also very important to distinguish between the case where $h_{j}$ are $N\times N$ blocks with $N>1$ and the scalar case, $N=1$, where $h_{j}$ are just complex numbers. In the latter case, Coburn’s lemma states that either $D(\text{ker}(H))=0$ or $D(\text{ker}(H^{\dagger}))=0$. I.e., in this case the sign of a non-zero winding number $\mathcal{I}$ tells us immediately whether $H$ has only right or only left zeroes and $|\mathcal{I}|$ gives the total number of such zeroes.
Example 1: To illustrate these points, we consider an example where $h(k)=1+2\text{e}^{-\text{i}k}$ which means that $\mathcal{I}=-\text{ind}(H)=-1$ and the Fourier coefficients are $h_{0}=1$ and $h_{1}=2$ with all the other ones being equal to zero. Coburn’s lemma then implies that $D(\text{ker}(H))=1$ and $D(\text{ker}(H^{\dagger}))=0$, i.e., $H$ has exactly one right zero mode. In this simple case, we can explicitly construct the right zero mode by considering $Hv=0$ with $v=\begin{pmatrix}v_{1}&v_{2}&\dots\end{pmatrix}^{T}$ and $v_{j}\in\mathbb{C}$. This leads to the recurrence relation for the vector coefficients $v_{j}+2v_{j+1}=0$. A normalized solution is given by $v_{j}=(-1)^{j-1}\sqrt{3}/2^{j}$ which means that the zero mode is exponentially localized.
@@ IV Finite systems and hidden zero modes
In a finite Hermitian system with non-trivial SPT order and system size $L$, zero modes can typically be easily identified as eigenvalues which scale as $|\lambda|\sim\exp(-L)$. This, however, is in general not the case in a non-Hermitian system. Here, zero modes can be hidden. This can be understood by considering the example discussed above. If we consider $H_{L}v=0$ for a finite matrix $H_{L}$ and vector $v=\begin{pmatrix}v_{1}&v_{2}&\dots&v_{L}\end{pmatrix}^{T}$ then the recurrence relation above is supplemented by the boundary condition $v_{L+1}=0$. In this case, the only solution is $v\equiv 0$ which is not a proper solution. The protected zero mode, which does exist in the thermodynamic limit, cannot be found by considering the eigenspectrum of a finite system. The only eigenstate in the finite case is $v=\begin{pmatrix}1&0&\dots 0\end{pmatrix}^{T}$ with eigenvalue $E=1$.
@@
In the example, the vector $v=\begin{pmatrix}v_{1}&v_{2}&\dots&v_{L}\end{pmatrix}^{T}$ with $v_{j}=(-1)^{j-1}\sqrt{3}/2^{j}$ is a hidden zero mode with $\lim_{L\to\infty}||H_{L}v||=\lim_{L\to\infty}|v_{L}|=0$.
@@
Example 2: This is best illustrated by another example. Consider the Hamiltonian in $k$-space given by
@@
Figure 1: (a) The Hamiltonian $H$ in Eq. (7) and (b) its adjoint $H^{\dagger}$; (c) its reflection $\tilde{H}$ and (d) its adjoint $\tilde{H}^{\dagger}$. The arrows indicate unidirectional hopping and the open circles the localized edge modes.
If we consider this Hamiltonian on a finite line then we find two exact zero modes which are localized at the right boundary. However, adding a small hopping in the opposite direction, one zero mode can be removed and the other one becomes a hidden zero mode.
@@ V Singular values and hidden zero modes
In the block case, $N>1$, both $D(\mbox{ker}(H))$ and $D(\mbox{ker}(H^{\dagger}))$ can be non-zero and the winding number $\mathcal{I}$ only fixes the difference. We know that $\tilde{\mathcal{I}}=-\mathcal{I}$ which implies that we can also write $K=D(\mbox{ker}(H^{\dagger}))+D(\mbox{ker}(\tilde{H}^{\dagger}))$ and therefore
@@
I.e., in the block case we only know that there are at least $|\mathcal{I}|\leq K$ singular values which will go to zero. Note that this is similar to the case of a Hermitian Chern insulator where the Chern number also only fixes the difference between left and right propagating modes but not their total number. The Chern number therefore also only provides a lower bound on the total number of edge modes.
@@
where we have used that $U$ is unitary. We thus find one of the main results of this letter: Every singular value $s_{n}$ which vanishes in the thermodynamic limit is directly connected to a (hidden) zero mode $v_{n}$ of the Hamiltonian $H_{L}$. The $K$-splitting theorem then connects this to the winding number. For a system with winding number $\mathcal{I}$ there are at least $|\mathcal{I}|$ exact or hidden zero modes. In the scalar case $N=1$, there are exactly $|\mathcal{I}|$ many which are protected by the winding number.
@@ VI Hidden zero modes in a sub-lattice symmetric multiband model
The case $\alpha=-1$ further exemplifies some of the findings in this work. Now $\mathcal{I}_{2}=-1$ if $x>0.9y$ and $\mathcal{I}_{2}=0$ otherwise meaning that the total winding $\mathcal{I}$ is either zero or $1$. Naively, we might expect that for $\mathcal{I}=0$ the topology is trivial and there is no skin effect. This, however, is incorrect. With sub-lattice symmetry being present, there are two topological invariants which can be either considered to be the windings of the blocks $\mathcal{I}_{1,2}$ or the sum and the difference of these two invariants [15, 32, 33]. The number of singular values $K$ is bounded from below by $K\geq|\mathcal{I}_{1}|+|\mathcal{I}_{2}|$. We thus expect to have either two or just one protected singular value consistent with the numerical results shown in Fig. 3. However, the single zero mode for the $\mathcal{I}_{1}=1$, $\mathcal{I}_{2}=0$ case is hidden while two eigenvalues which go to zero are present for the $\mathcal{I}_{1}=1$, $\mathcal{I}_{2}=-1$ case. The latter can be explained by noting that the system for $\mathcal{I}_{1}=-\mathcal{I}_{2}$ can be adiabatically deformed to a Hermitian one. The two zero eigenmodes then follow from the standard bulk-boundary correspondence.